# Iyama–Solberg correspondence for exact dg categories

### 04 Jan 2024 v1

@@ 3.13. Connection to higher Auslander–Solberg correspondence for exact categories
In the rest of this section, we assume $d\geq 1$ and compare our correspondence with the correspondence obtained by Grevstad in [11]. By combining Proposition 3.18 and Lemma 3.19, we see that we get a generalization of the correspondence [11, Theorem 3.8] in the case of weakly idempotent complete exact categories with enough injectives.
Definition 3.14 ([7, Proposition 1.7]).
@@
is called acyclic or exact if each $d^{i}$ is admissible and $\mathrm{ker}(d^{i+1})=\mathrm{im}(d^{i})$. Let $\mathcal{M}$ be a full subcategory of $\mathcal{E}$. We write $\mathrm{mod}_{adm}(\mathcal{M})$ [9, Definition 3.1 (2)] for the full subcategory of $\mathrm{Mod}\,\mathcal{M}$ consisting of those functors $F$ that admit a projective presentation
@@
Put $\mathrm{cogen}(\mathcal{X})=\{E\in\mathcal{E}\mid\exists\text{ inflation }E\rightarrowtail X\text{ with }X\in\mathcal{X}\}$. A pair $(\mathcal{T},\mathcal{F})$ of full subcategories of $\mathcal{E}$ is a torsion pair [12, Definition 2.17] if
@@
The above bijection is a restriction of the bijection in [11, Theorem 3.8]. More precisely, in Proposition 6.3 in loc. cit, it is shown that in the bijection of Theorem 3.8 in loc. cit, the exact categories $\mathcal{E}$ has enough injectives if and only if for the corresponding exact category $\mathcal{E}^{\prime}$ we have $\mathrm{domdim}(\mathcal{E}^{\prime})\geq 1$. But since $\mathcal{M}$ is $d$-rigid, we have $\mathrm{domdim}(\mathcal{E}^{\prime})\geq d+1$ in this case. By [12, Lemma 4.17 (2)], if $\mathcal{E}^{\prime}$ has enough projectives and if $\mathcal{P}$ is admissibly covariantly finite and $\mathrm{domdim}(\mathcal{E}^{\prime})\geq 1$, then $({{}^{\perp}\mathcal{P}},\mathrm{cogen}(\mathcal{P}))$ is a torsion pair. By [12, Lemma 4.15], if $\mathcal{E}^{\prime}$ has enough projectives and $\mathrm{domdim}(\mathcal{E}^{\prime})\geq 1$, then ${{}^{\perp}\mathcal{P}}$ is closed under admissible subobjects. Since $\mathrm{domdim}(\mathcal{E}^{\prime})\geq d+1$, a similar proof of [12, Lemma 4.14] shows that $\mathrm{Ext}^{1\sim d}_{\mathcal{E}}({{}^{\perp}\mathcal{P}},\mathcal{P})=0$. Therefore the exact categories appearing in (2) of the above bijection are partial $d$-minimal Auslander–Gorenstein in the sense of [11, Definition 3.4].

### 09 Feb 2024 v2

@@ 3.13. Connection to higher Auslander–Solberg correspondence for exact categories
In the rest of this section, we assume $d\geq 1$ and compare our correspondence with the correspondence obtained by Grevstad in [13]. By combining Proposition 3.18 and Lemma 3.19, we see that we get a generalization of the correspondence [13, Theorem 3.8] in the case of weakly idempotent complete exact categories with enough injectives.
Definition 3.14 ([8, Proposition 1.7]).
@@
is called acyclic or exact if each $d^{i}$ is admissible and $\mathrm{ker}(d^{i+1})=\mathrm{im}(d^{i})$. Let $\mathcal{M}$ be a full subcategory of $\mathcal{E}$. We write $\mathrm{mod}_{adm}(\mathcal{M})$ [10, Definition 3.1 (2)] for the full subcategory of $\mathrm{Mod}\,\mathcal{M}$ consisting of those functors $F$ that admit a projective presentation
@@
Put $\mathrm{cogen}(\mathcal{X})=\{E\in\mathcal{E}\mid\exists\text{ inflation }E\rightarrowtail X\text{ with }X\in\mathcal{X}\}$. A pair $(\mathcal{T},\mathcal{F})$ of full subcategories of $\mathcal{E}$ is a torsion pair [14, Definition 2.17] if
@@
The above bijection is a restriction of the bijection in [13, Theorem 3.8]. More precisely, in Proposition 6.3 in loc. cit, it is shown that in the bijection of Theorem 3.8 in loc. cit, the exact categories $\mathcal{E}$ has enough injectives if and only if for the corresponding exact category $\mathcal{E}^{\prime}$ we have $\mathrm{domdim}(\mathcal{E}^{\prime})\geq 1$. But since $\mathcal{M}$ is $d$-rigid, we have $\mathrm{domdim}(\mathcal{E}^{\prime})\geq d+1$ in this case. By [14, Lemma 4.17 (2)], if $\mathcal{E}^{\prime}$ has enough projectives and if $\mathcal{P}$ is admissibly covariantly finite and $\mathrm{domdim}(\mathcal{E}^{\prime})\geq 1$, then $({{}^{\perp}\mathcal{P}},\mathrm{cogen}(\mathcal{P}))$ is a torsion pair. By [14, Lemma 4.15], if $\mathcal{E}^{\prime}$ has enough projectives and $\mathrm{domdim}(\mathcal{E}^{\prime})\geq 1$, then ${{}^{\perp}\mathcal{P}}$ is closed under admissible subobjects. Since $\mathrm{domdim}(\mathcal{E}^{\prime})\geq d+1$, a similar proof of [14, Lemma 4.14] shows that $\mathrm{Ext}^{1\sim d}_{\mathcal{E}}({{}^{\perp}\mathcal{P}},\mathcal{P})=0$. Therefore the exact categories appearing in (2) of the above bijection are partial $d$-minimal Auslander–Gorenstein in the sense of [13, Definition 3.4].